# Quantum mixing on large Schreier graphs

### v1

@@ 4.3 Norm estimation with RD property
$\|Q\|_{\rm op}$ $=\|\sum_{g}\bar{f}_{1}(p)_{t^{-1}gt}f_{2}(p)_{g}\lambda(g)\|_{\rm op}$
$\leq\|\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}\gamma^{z_{2}}_{g}\lambda(g)\|_{\rm op}+\|\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}a_{2,g}\lambda(g)\|_{\rm op}+\|\sum_{g}\bar{a}_{1,t^{-1}gt}f_{2}(p)_{g}\lambda(g)\|_{\rm op}.$
We next use $\|\sum_{g\in\Gamma}a_{g}\lambda(g)\|_{\rm op}\leq\|a\|_{1}$ for $a\in\mathbb{C}[\Gamma]$ and $\|a_{i}\|_{2}=\|a_{i}\delta_{e}\|_{2}\leq\|a_{i}\|_{\rm op}\leq 4e^{-1/\epsilon}$, we find for some $C>0$,
$\|Q\|_{\rm op}$ $\leq\|\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}\gamma^{z_{2}}_{g}\lambda(g)\|_{\rm op}+\sum_{g}|\gamma^{z_{1}}_{1,gt}a_{2,tg}|+\sum_{g}|a_{1,gt}f_{2}(p)_{tg}|$
$\leq\|\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}\gamma^{z_{2}}_{g}\lambda(g)\|_{\rm op}+4e^{-1/\epsilon}(\|\gamma^{z_{1}}\|_{2}+\|f_{2}(p)\|_{2})$
$\leq\|\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}\gamma^{z_{2}}_{g}\lambda(g)\|_{\rm op}+C\eta^{3/2},$
@@
We can now use RD decay property to evaluate the norm of $\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}\gamma^{z_{2}}_{g}\lambda(g)$. If $C^{\prime}_{1}>2C_{1}+1$, we get using Lemma B.1
$\|\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}\gamma^{z_{2}}_{g}\lambda(g)\|_{\rm op}\leq C\sqrt{\sum_{g}(|g|+1)^{C^{\prime}_{1}}|\gamma^{z_{1}}_{t^{-1}gt}|^{2}|\gamma^{z_{2}}_{g}|^{2}}.$
@@
$\|\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}\gamma^{z_{2}}_{g}\lambda(g)\|_{\rm op}$ $\leq C\sqrt{\sum_{g}(|gt|+1)^{C^{\prime}_{1}}|\gamma^{z_{1}}_{gt}|^{2}|\gamma^{z_{2}}_{gt}|^{2}}$
@@
In this subsection, we prove Theorem 1.7, i.e. quantum mixing under strong convergence. We build upon Sections 4.1 and 4.3. Recall that $Q_{N}=\rho_{N}(q)$ and $q$ depends on $(\eta,E_{1},E_{2})$. It is enough to prove that (4.9) holds for all $E_{1},E_{2}\in I$. We consider a subset $\mathcal{E}=\mathcal{E}_{\eta}$ of $C/\eta$ points in $I$ such that all $E\in I$ are within distance $\eta$ to $\mathcal{E}$. Then by Lemma 2.2, it suffices to prove that
$\lim_{\eta\to 0}\max_{E_{1},E_{2}\in\mathcal{E}_{\eta}}\limsup_{N\to\infty}\frac{\langle k_{N},Q_{N}k_{N}\rangle}{N\eta}=0.$ (4.12)
@@
Since $\|k_{N}\|_{\infty}\leq 1$, we get $\|k_{N}\|_{2}\leq\sqrt{N}$ and thus
@@
Also, we can find $N_{0}=N_{0}(\eta)$ such that (4.13) holds for all $N\geq N_{0}$ and all $E_{1},E_{2}\in\mathcal{E}_{\eta}$. From (4.11), we deduce that (4.12) holds. It concludes the proof. ∎
@@
More precisely, since $\rho_{N}$ converges in distribution toward $\lambda$, we can find thanks to a diagonal extraction process a sequence $r_{N}\to\infty$ such that $|{\rm Bad}_{S}(r_{N})|/N\to 0$, where as usual, for integer $r\geq 0$, ${\rm Bad}_{S}(r)=\mathrm{Fix}^{*}_{N}(B_{S}(2r))$. We may further find integer sequences $n_{N}\to\infty$, $\ell_{N}\to\infty$ such that
@@
(for instance $n_{N}=\ell_{N}=\lfloor\sqrt{r_{N}}\rfloor$).
Next, we may find a sequence $\eta_{N}\to 0$ such that
$n_{N}\geq\frac{c}{\eta_{N}\epsilon_{N}}\;,\quad\lim_{N\to\infty}\frac{|{\rm Bad}_{S}(n_{N}\ell_{N})|}{N\eta_{N}^{2}}=0\quad\hbox{ and }\quad\lim_{N\to\infty}\eta_{N}^{2}4^{\ell_{N}}=+\infty.$ (4.14)
with $\epsilon_{N}=1/(\ln(\max(1/\eta_{N},4)))$ as in (4.3) and $c>0$ as in Lemma 2.3.
As usual, let $q=q_{t,E_{1},E_{2},\eta_{N}}\in\mathbb{C}[\Gamma]$ as defined in (4.7), $Q_{N}=\rho_{N}(q)$ and $Q=\lambda(q)$ (they depend implicitly on $(t,E_{1},E_{2},\eta_{N})$). Let $\mathcal{E}_{\eta_{N}}\subset I$ be the subset of $C/\eta_{N}$ points in $I$ defined above (4.12). To prove Theorem 1.5, it is enough to prove that there exists a vector subspace $H_{N}\subset\mathbb{C}^{N}$ of dimension $\mathrm{dim}(H_{N})=o(N)$ such that for any $k_{N}\in H_{N}^{\perp}$ with $\|k_{N}\|_{\infty}\leq 1$, we have
$\lim_{N\to\infty}\max_{E_{1},E_{2}\in\mathcal{E}_{\eta_{N}}}\frac{\langle k_{N},Q_{N}k_{N}\rangle}{N\eta_{N}}=0.$ (4.15)
@@
$\leq|{\rm Bad}_{S}(n_{N}\ell_{N})|+N2^{-2\ell_{N}}.$
In particular, for each $E_{1},E_{2}\in I$, the vector space $H_{N,E_{1},E_{2}}$ spanned by eigenvectors of $Q_{N}=\rho_{N}(q_{t,E_{1},E_{2},\eta_{N}})$ associated with eigenvalues larger than $2\|Q\|_{\rm op}$ has dimension at most $|{\rm Bad}_{S}(n_{N}\ell_{N})|+N2^{-2\ell_{N}}$. Let $H_{N}=\sum_{E_{1},E_{2}\in\mathcal{E}_{\eta_{N}}}H_{N,E_{1},E_{2}}$. This vector space has dimension at most
$|\mathcal{E}_{\eta_{N}}|^{2}\left(|{\rm Bad}_{S}(n_{N}\ell_{N})|+N2^{-2\ell_{N}}\right)=o(N),$
where we have used $|\mathcal{E}_{\eta_{N}}|\leq C/\eta_{N}$ and (4.14).
On the other hand, if $k_{N}\in H_{N}^{\perp}$, then for any $E_{1},E_{2}\in\mathcal{E}_{\eta_{N}}$, we have
@@
Finally, we use (4.11) and this concludes the proofs of (4.15) and of Theorem 1.5.
@@
Since $\mathbb{E}a_{N}(x)$ does not depend on $x$, by linearity we may assume without loss of generality that $\mathbb{E}a_{N}(x)=0$. Also, Hoeffding’s inequality (see [19, Theorem 2.8]) and Borel-Cantelli lemma imply that almost-surely
@@
$|F(a)-F(b)|\leq\frac{1}{N}\sum_{x=1}^{N}2\cdot\mathbf{1}(a_{x}\neq b_{x}).$
@@
$\mathbb{P}\bigg(\Big|\frac{1}{N}\sum_{x\in C_{N}}a_{N}(x)\bar{a}_{N}(g.x)\Big|\geq t\bigg)\leq 2e^{-\frac{Nt^{2}}{2}}.$
@@ 5.1 Setup
In this section, as announced in SubSection 1.5, we state an extension of our main theorems to the case of finite coverings. This is an important extension in view of applications, notably to coverings of graphs and product of graphs, see Examples 5.1-5.2 below.
@@
The operators (5.1) and (5.2) have a probabilistic interpretation if all coefficients $p_{g}(i,j)$ are non-negative and $\sum_{g\in\Gamma}p_{g}$ is a stochastic matrix. Indeed then $\lambda(p)$ is the transition matrix of a random walk on $\left[r\right]\times\Gamma$ where the probability to jump from $(i,g)$ to $(j,h)$ is $p_{gh^{-1}}(i,j)$. We can think as an element of $\left[r\right]$ as a color, as $\left[r\right]\times\{g\}$ as the fiber above $g\in\Gamma$ and as $\{i\}\times\Gamma$ as the copy of $\Gamma$ with color $i$. The same interpretation holds for $\rho_{N}(p)$.
@@
The adjacency matrix of an $N$-covering of a base graph $H=(V,E)$ with vertex set $V=[r]$ can also be put in the form (5.2). If $d=|E|$ denotes the number of undirected edges of $H$, let $\Gamma=F_{d}$ be the free group with $d$ free generators and their $d$ inverses denoted by $(g_{1},\ldots,g_{2d})$, $g_{i^{*}}=g_{i}^{-1}$ for some involution $i\mapsto i^{*}$ on $\left[2d\right]$. Each generator $g_{i}$ is associated to a directed edge $g_{i}=(u_{i},v_{i})$ of $H$, and $g_{i}^{-1}=g_{i^{*}}=(v_{i},u_{i})$. We consider the action of $\Gamma$ on $[N]$ defined by $\rho_{N}(g_{i})(x)=\sigma_{i}(x)$ where $(\sigma_{1},\ldots,\sigma_{d})$ are any permutation matrices such that $\sigma_{i}^{-1}=\sigma_{i^{*}}$. Then an $N$-covering is a graph whose adjacency matrix can be written as $\rho_{N}(1_{E})=\sum_{i}E_{u_{i},v_{i}}\otimes\rho_{N}(g_{i})$ where $E_{u_{i},v_{i}}\in M_{r}(\mathbb{R})$ is the matrix whose only non-zero element is $(u_{i},v_{i})$ (equal to $1$) and $1_{E}=\sum_{i}E_{u_{i}v_{i}}g_{i}\in M_{r}(\mathbb{C})[F_{d}]$. The vertex $(u,x)\in\left[r\right]\times\left[N\right]$ is connected to vertex $(v,y)$ through the $i$-th directed edge if $(u_{i},v_{i})=(u,v)$ and $y=\sigma_{i}(x)$. The operator $\lambda(1_{E})=\sum_{i}E_{u_{i},v_{i}}\otimes\lambda(g_{i})$ is the adjacency operator of $r$ disjoint copies of the universal covering tree of $H$. By replacing the free group $F_{d}$ by $\mathbb{Z}^{d}$, we obtain maximal Abelian coverings (see [13] for the more general notion of $(\Gamma,S)$-covering graphs).

### v2

@@ 4.3 Norm estimation with RD property
$\|Q\|_{\rm op}$ $=\bigg\|\sum_{g}\bar{f}_{1}(p)_{t^{-1}gt}f_{2}(p)_{g}\lambda(g)\bigg\|_{\rm op}$
$\leq\bigg\|\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}\gamma^{z_{2}}_{g}\lambda(g)\bigg\|_{\rm op}+\bigg\|\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}a_{2,g}\lambda(g)\bigg\|_{\rm op}+\bigg\|\sum_{g}\bar{a}_{1,t^{-1}gt}f_{2}(p)_{g}\lambda(g)\bigg\|_{\rm op}.$
We next use $\|\sum_{g\in\Gamma}a_{g}\lambda(g)\|_{\rm op}\leq\|a\|_{1}$ for $a\in\mathbb{C}[\Gamma]$ and $\|a_{i}\|_{2}=\|a_{i}\delta_{e}\|\leq\|a_{i}\|_{\rm op}\leq 4e^{-1/\epsilon}$. This gives, for some $C>0$,
$\|Q\|_{\rm op}$ $\leq\bigg\|\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}\gamma^{z_{2}}_{g}\lambda(g)\bigg\|_{\rm op}+\sum_{g}|\gamma^{z_{1}}_{1,gt}a_{2,tg}|+\sum_{g}|a_{1,gt}f_{2}(p)_{tg}|$
$\leq\bigg\|\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}\gamma^{z_{2}}_{g}\lambda(g)\bigg\|_{\rm op}+4e^{-1/\epsilon}(\|\gamma^{z_{1}}\|_{2}+\|f_{2}(p)\|_{2})$
$\leq\bigg\|\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}\gamma^{z_{2}}_{g}\lambda(g)\bigg\|_{\rm op}+C\eta^{3/2},$
@@
We can now use the RD property to evaluate the norm of $\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}\gamma^{z_{2}}_{g}\lambda(g)$. If $C^{\prime}_{1}>2C_{1}+1$, we get using Lemma B.1,
$\bigg\|\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}\gamma^{z_{2}}_{g}\lambda(g)\bigg\|_{\rm op}\leq C\sqrt{\sum_{g}(|g|+1)^{C^{\prime}_{1}}|\gamma^{z_{1}}_{t^{-1}gt}|^{2}|\gamma^{z_{2}}_{g}|^{2}}.$
@@
$\bigg\|\sum_{g}\bar{\gamma}^{z_{1}}_{t^{-1}gt}\gamma^{z_{2}}_{g}\lambda(g)\bigg\|_{\rm op}$ $\leq C\sqrt{\sum_{g}(|gt|+1)^{C^{\prime}_{1}}|\gamma^{z_{1}}_{gt}|^{2}|\gamma^{z_{2}}_{tg}|^{2}}$
@@
In this subsection, we prove Theorem 1.7, i.e. quantum mixing under strong convergence. We build upon Sections 4.1 and 4.3. Recall that $Q_{N}=\rho_{N}(q)$ and $q$ depends on $(\eta,E_{1},E_{2})$. It is enough to prove that (4.9) holds for all $E_{1},E_{2}\in I$. We consider a subset $\mathcal{E}=\mathcal{E}_{\eta}$ of $C/\eta$ points in $I$ such that all $E\in I$ are within distance $\eta$ to $\mathcal{E}$. Then, by the discussion above (4.9), it suffices to prove that
$\lim_{\eta\to 0}\limsup_{N\to\infty}\max_{E_{1},E_{2}\in\mathcal{E}_{\eta}}\frac{\langle k_{N},Q_{N}k_{N}\rangle}{N\eta}=0.$ (4.12)
@@
Using that $\|k_{N}\|_{\infty}\leq 1$, we get $\|k_{N}\|_{2}\leq\sqrt{N}$ and thus
@@
We may choose a large $N_{0}=N_{0}(\eta)$ such that (4.13) holds for all $N\geq N_{0}$ and all $E_{1},E_{2}\in\mathcal{E}_{\eta}$. From (4.11), we deduce that (4.12) holds. It concludes the proof. ∎
@@
Since $\rho_{N}$ converges in distribution to $\lambda$, we can find thanks to a diagonal extraction process a sequence $r_{N}\to\infty$ such that $|{\rm Bad}_{S}(r_{N})|/N\to 0$, where as usual, for integer $r\geq 0$, ${\rm Bad}_{S}(r)=\mathrm{Fix}^{*}_{N}(B_{S}(2r))$. We may further find integer sequences $n_{N}\to\infty$, $\ell_{N}\to\infty$ such that
@@
(for instance $n_{N}=\ell_{N}=\lfloor\sqrt{r_{N}}\rfloor$). Let
$f_{N}=\frac{N}{|{\rm Bad}_{S}(n_{N}\ell_{N})|+N4^{-\ell_{N}}}.$ (4.14)
Then $f_{N}\rightarrow+\infty$.
Fix $m\geq 2$, and set $\eta_{m}=2^{-m}$. As in (4.3), let $\epsilon_{m}=1/\log(1/\eta_{m})$ and $N_{m}$ large enough so that $\eta_{m}\epsilon_{m}\geq c/n_{N}$ for any $N\geq N_{m}$, where $c>0$ is the numerical constant of Lemma 2.3. As usual, let $q=q_{t,E_{1},E_{2},\eta_{m}}\in\mathbb{C}[\Gamma]$ as defined in (4.7), $Q_{N}=\rho_{N}(q)$ and $Q=\lambda(q)$ (they depend implicitly on $(t,E_{1},E_{2},\eta_{m})$). Let $\mathcal{E}_{\eta_{m}}\subset I$ be the subset of $C/\eta_{m}$ points in $I$ defined above (4.12). To prove Theorem 1.5, it is enough to prove that there exists a vector subspace $H_{N}\subset\mathbb{C}^{N}$ of dimension $\mathrm{dim}(H_{N})=o(N)$ (note that $H_{N}$ depends on $N$ but not on $m$) such that for any $k_{N}\in H_{N}^{\perp}$ with $\|k_{N}\|_{\infty}\leq 1$, we have
$\lim_{m\rightarrow+\infty}\lim_{N\to\infty}\max_{E_{1},E_{2}\in\mathcal{E}_{\eta_{m}}}\frac{\langle k_{N},Q_{N}k_{N}\rangle}{N\eta_{m}}=0.$ (4.15)
Here we note that the definition (1.11) of small scale quantum mixing is written with the limit $\lim_{\eta\rightarrow 0}$, but it is actually equivalent to consider the limit along the subsequence $\eta_{m}\rightarrow 0$. Indeed, the quantity $\sum_{\alpha\in\Lambda_{J_{E_{1}}^{\eta}}}\sum_{\beta\in\Lambda_{J_{E_{2}}^{\eta}}}\left|\langle\varphi_{\beta},K_{N}\varphi_{\alpha}\rangle-\langle\varphi_{\beta},\langle K_{N}\rangle\varphi_{\alpha}\rangle\right|^{2}$ is non-decreasing in $\eta$, and the quantity $\frac{1}{|\Lambda_{J_{E_{1}}^{\eta}}|}$ is comparable to $N\eta$ owing to Lemma 3.2, hence it varies at most by a constant factor on dyadic intervals $\eta_{m},\eta_{m-1}]$. Thus we do not lose generality by focusing on the case where $\eta$ belongs to the sequence $(\eta_{m})$.
@@
$\leq|{\rm Bad}_{S}(n_{N}\ell_{N})|+N4^{-\ell_{N}}.$
In particular, for each $E_{1},E_{2}\in I$, the vector space $H_{N,E_{1},E_{2},m}$ spanned by eigenvectors of $Q_{N}=\rho_{N}(q_{t,E_{1},E_{2},\eta_{m}})$ associated with eigenvalues larger than $2\|Q\|_{\rm op}$ has dimension at most $|{\rm Bad}_{S}(n_{N}\ell_{N})|+N4^{-\ell_{N}}$. Consider the sum of vector spaces $H_{N}^{(m)}=\sum_{E_{1},E_{2}\in\mathcal{E}_{\eta_{m}}}H_{N,E_{1},E_{2},\eta_{m}}$. Since $|\mathcal{E}_{\eta_{m}}|\leq C/\eta_{m}$,
$\dim(H_{N}^{(m)})\leq|\mathcal{E}_{\eta_{m}}|^{2}\left(|{\rm Bad}_{S}(n_{N}\ell_{N})|+N4^{-\ell_{N}}\right)\leq C^{2}4^{m}\frac{N}{f_{N}}.$ (4.16)
For $N\in\mathbb{N}$, we finally form the sum of vector spaces
$H_{N}=\sum_{1\leq m\leq\alpha_{N}}H_{N}^{(m)}$
where $\alpha_{N}\geq 0$ is the largest integer such that
$\sum_{1\leq m\leq\alpha_{N}}\dim(H_{N}^{(m)})\leq\frac{N}{\sqrt{f_{N}}}.$ (4.17)
If $\alpha_{N}=0$, then we let $H_{N}=\emptyset$. By definition, $\dim(H_{N})\leq\frac{N}{\sqrt{f_{N}}}=o(N)$. Moreover, $\alpha_{N}\rightarrow+\infty$ (due to (4.16) and the fact that $f_{N}\rightarrow+\infty$), and therefore for any $m\in\mathbb{N}$, there exists $N_{m}^{\prime}\geq N_{m}$ such that $H_{N}^{(m)}$ appears in $H_{N}$ for any $N\geq N_{m}$.
Finally, for $m\in\mathbb{N}$, for any $E_{1},E_{2}\in\mathcal{E}_{\eta_{m}}$, if $N\geq N_{m}^{\prime}$ and $k_{N}\in H_{N}^{\perp}$, then $k_{N}\in(H_{N}^{(m)})^{\perp}$, and therefore
@@
Using (4.11), this concludes the proofs of (4.15) and of Theorem 1.5.
@@
We prove here that random observables are uncorrelated. Since $\mathbb{E}a_{N}(x)$ does not depend on $x$, by linearity we may assume without loss of generality that $\mathbb{E}a_{N}(x)=0$. Also, Hoeffding’s inequality (see [19, Theorem 2.8]) and Borel-Cantelli lemma imply that almost-surely
@@
$|F(a)-F(b)|\leq\frac{1}{N}\sum_{x=1}^{N}4\cdot\mathbf{1}(a_{x}\neq b_{x}).$
@@
$\mathbb{P}\bigg(\Big|\frac{1}{N}\sum_{x\in C_{N}}a_{N}(x)\bar{a}_{N}(g.x)\Big|\geq t\bigg)\leq 2e^{-\frac{Nt^{2}}{8}}.$
@@ 5.1 Setup
In this section, as announced in Section 1.5, we state an extension of our main theorems to the case of finite coverings. This is an important extension in view of applications, notably to coverings of graphs and product of graphs, see Examples 5.1-5.2 below.
@@
The operators (5.1) and (5.2) have a probabilistic interpretation if all coefficients $p_{g}(i,j)$ are non-negative and $\sum_{g\in\Gamma}p_{g}$ is a stochastic matrix. Indeed then $\lambda(p)$ is the transition matrix of a random walk on $\left[r\right]\times\Gamma$ where the probability to jump from $(i,g)$ to $(j,h)$ is $p_{gh^{-1}}(i,j)$. We can think of an element of $\left[r\right]$ as a color, of $\left[r\right]\times\{g\}$ as the fiber above $g\in\Gamma$ and of $\{i\}\times\Gamma$ as the copy of $\Gamma$ with color $i$. The same interpretation holds for $\rho_{N}(p)$.
@@
The adjacency matrix of an $N$-covering of a base graph $H=(V,E)$ with vertex set $V=[r]$ can also be put in the form (5.2). If $d=|E|$ denotes the number of undirected edges of $H$, let $\Gamma=F_{d}$ be the free group with $d$ free generators and their $d$ inverses denoted by $(g_{1},\ldots,g_{2d})$, $g_{i^{*}}=g_{i}^{-1}$ for some involution $i\mapsto i^{*}$ on $\left[2d\right]$. Each generator $g_{i}$ is associated to a directed edge $g_{i}=(u_{i},v_{i})$ of $H$, and $g_{i}^{-1}=g_{i^{*}}=(v_{i},u_{i})$. We consider the action of $\Gamma$ on $[N]$ defined by $\rho_{N}(g_{i})(x)=\sigma_{i}(x)$ where $(\sigma_{1},\ldots,\sigma_{d})$ are any permutation matrices such that $\sigma_{i}^{-1}=\sigma_{i^{*}}$. Then an $N$-covering is a graph whose adjacency matrix can be written as $\rho_{N}(1_{E})=\sum_{i}E_{v_{i}u_{i}}\otimes\rho_{N}(g_{i})$ where $E_{v_{i}u_{i}}\in M_{r}(\mathbb{R})$ is the matrix whose only non-zero element is $(u_{i},v_{i})$ (equal to $1$) and $1_{E}=\sum_{i}E_{v_{i}u_{i}}g_{i}\in M_{r}(\mathbb{C})[F_{d}]$. The vertex $(u,x)\in\left[r\right]\times\left[N\right]$ is connected to vertex $(v,y)$ through the $i$-th directed edge if $(u_{i},v_{i})=(u,v)$ and $y=\sigma_{i}(x)$. The operator $\lambda(1_{E})=\sum_{i}E_{v_{i}u_{i}}\otimes\lambda(g_{i})$ is the adjacency operator of $r$ disjoint copies of the universal covering tree of $H$. Note that $H$ can be a multigraph. By replacing the free group $F_{d}$ by $\mathbb{Z}^{d}$, we obtain maximal Abelian coverings (see [13] for the more general notion of $(\Gamma,S)$-covering graphs).